OpenStudy (anonymous):

Question 2 (Essay Worth 10 points) (07.02 MC) An equation is shown below: 3(2x – 7) = 3 Part A: How many solutions does this equation have? (4 points) Part B: What are the solutions to this equation? Show your work. (6 points)

OpenStudy (anonymous):

Question 3 (Essay Worth 10 points) (07.03 MC) An equation is shown below: 9(3x – 16) + 15 = 6x – 24 Part A: Write the steps you will use to solve the equation, and explain each step. (6 points) Part B: What value of x makes the equation true? (4 points)

OpenStudy (anonymous):

Question 4 (Essay Worth 10 points) (07.05 MC) An equation is shown below: 4x + 2(x – 3) = 4x + 2x – 11 Part A: Solve the equation and write the number of solutions. Show all the steps. (6 points) Part B: Name one property you used to solve this equation. (4 points)

OpenStudy (anonymous):

medal and and fan thingy

OpenStudy (anonymous):

q1. the whole equation shows x=4 so 1 solution

OpenStudy (anonymous):

your first question i mean

OpenStudy (anonymous):

3*2x=6x

OpenStudy (anonymous):

3*-7=-21

OpenStudy (anonymous):

so 6x-21=3

OpenStudy (anonymous):

then you add 21 on both sides

OpenStudy (anonymous):

then u get 6x=24 then u divide both sides by 6 and ur answer is 4

OpenStudy (anonymous):

that was for the first one

OpenStudy (anonymous):

i'm afraid i have to go im very sorry but i have to do 56 questions as homework just use distribution and remember PEMDAS

OpenStudy (anonymous):

thank you @GTA_Hunter35

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

:D